跳到论文正文
W Papers独立论文阅读器 arXiv:2212.14555 论文正文 计算阅读时间 · 0%
引用这篇论文根据当前论文元数据生成,正式投稿前建议核对。

    
  
License: CC BY 4.0
arXiv:2212.14555v3 [stat.ML] 28 May 2023

Relative Probability on Finite Outcome Spaces
A Systematic Examination of its Axiomatization, Properties, and Applications

Max Sklar Affiliation: Local Maximum Labs Affiliation: December 30, 2022
Abstract

This work proposes a view of probability as a relative measure rather than an absolute one. To demonstrate this concept, we focus on finite outcome spaces and develop three fundamental axioms that establish requirements for relative probability functions. We then provide a library of examples of these functions and a system for composing them. Additionally, we discuss a relative version of Bayesian inference and its digital implementation. Finally, we prove the topological closure of the relative probability space, highlighting its ability to preserve information under limits.

1 Introduction

The foundations of probability theory are still very much open to explore!

Since Kolmogorov published the standard axioms for probability[1] in 1933, there have been calls to alter them. In “Kolmogorov’s Axiomatisation and Its Discontents”, Lyon[2] examines arguments made in favor and against these alterations. One area of “discontentment” concerns conditional probability. We often want to identify the probability of event A given event B, or P⁡(A|B)P(A|B), even when BB has probability zero11 1 Note that zero probability events can indeed occur, particularly when given a continuous distribution. November[3] gave a more recent philosophical treatment of this phenomenon..

We are out of luck with the Kolmogorov model, which defines P⁡(A|B)P(A|B) as the ratio P⁡(A∩B)P⁡(B)\frac{P(A\cap B)}{P(B)}. When P⁡(B)=0P(B)=0, the indeterminate form 00\frac{0}{0} appears leaving the conditional probability undefined. In fact, this happens whenever one wishes to compare two events that have probability zero.

Undeterred by the petty obstacle of the indeterminate form, mathematicians and engineers refer to relative probabilities of this type all the time. For example, if we consider a probability distribution over [0,1][0,1] given by f⁡(x)=2​xf(x)=2x, we know that the value of f⁡(12)=1f(\frac{1}{2})=1 is twice as much as f⁡(14)=12f(\frac{1}{4})=\frac{1}{2}. In a sense, we believe that the outcome 12\frac{1}{2} is twice as likely as outcome 14\frac{1}{4} even though we are only talking about probability density.

Hajek[4] (citing Borel) gives a much more compelling example: if a random point on the Earth is selected, what is the probability that it is in the eastern hemisphere given that it is on the equator? It seems that one should not hesitate to answer one half. And yet the equator, being a mere 1-dimensional object, has probability zero compared to the rest of the globe.

To address the discrepancy between theory and practice, we propose a non-standard model of probability. This model is built on the relationships between outcomes and events rather than their absolute likelihoods. By focusing on these relationships, this improves on the Kolmogorov model by solving the conditional probability question. As additional benefits, it fits nicely with most distribution sampling techniques and establishes probability functions as categories.

1.1 Previous Work and Goals

Leading probability theorists in the twentieth century developed axiomatic systems for conditional probability, notably Renyi[5]. Later, Kohlberg and Reny[6] introduced the idea of relative probability and applied it to game theory. Heinemann[7] further developed the idea into relative probability measures along with their axioms and definitions, and found applications in Bayesian inference and economics[8].

This work is focused on the systematization of these concepts and will expand on them in several ways.

First, we will construct a theory of relative probability on finite outcome spaces. By omitting infinite outcome spaces, we temporarily set aside the need to account for measurable sets and countable additivity22 2 In the textbook Invitation to Discrete Mathematics[9], Matoušek et al. make the same simplification in order to illustrate concepts while at the same time writing By restricting ourselves to finite probability spaces we have simplified the situation considerably… A true probability theorist would probably say that we have excluded everything interesting. . Even with this vast simplification there is much to be learned. The fundamental definitions for relative probability will be more clearly constructed without the distractions that are introduced by continuous space.

Our focus on probability at the outcome level will lead us to separate out three fundamental axioms of relative probability. These fundamental axioms, acting on outcomes only, will be distinct from those related to summation and measurability. They will establish relative probability as a thin category on outcomes.

Second, we will categorize the various patterns and states that arise when dealing with relative probability, including total comparability, possibility classes, and anchor outcomes.

Third, we focus on the computational properties of Bayesian inference. The relative probability function will simplify the formulas for some distributions in the Bayesian framework. We introduce the indeterminate wildcard value, which will inevitably arise when relative probability is inferred on messy, real world data. This discussion ends with how these concepts can be applied to code and data structures.

Finally, we prove the ability of relative probability functions to retain information when taking limits, or in other words their topological closure.

Ultimately, practitioners will find these features of relative probability attractive. Its sphere of application could be expanded beyond game theory and theoretical probability into fields like machine learning. This paper provides a foundational analysis that future researchers can use to expand that sphere.

2 Preliminaries

2.1 Magnitude Space

Definition 2.1.

The magnitude space 𝕄\mathbb{M} is the set of all positive real numbers along with 00 and ∞\infty.

𝕄=[0,+∞]\mathbb{M}=[0,+\infty]

Magnitudes correspond with our intuition of size. The value of infinity is a limit element, larger that all of the other magnitudes. It endows the magnitude space with several important properties:

  1. 1.

    Compactness: Sequences that go off to infinity still have a limit (at ∞\infty).

  2. 2.

    Symmetry around ratios: When we compare the probability of two events, we get their odds. If the odds are 0, then we are comparing an event with probability 0 to an event with probability >0>0. We should be able to reverse this comparison, and say there are infinite odds when an event with probability >0>0 is compared to an event with probability 0. It is also common to define the odds of a single event as the odds of that event against its converse. In this case, ∞\infty corresponds to events that are certain.

  3. 3.

    The infinite element is introduced in measure theory because many mathematical systems (real and natural numbers for example) contain sets of infinite measure.

We set 0−1=∞0^{-1}=\infty and ∞−1=0\infty^{-1}=0, even though the product 0⋅∞0\cdot\infty is indeterminate.

2.2 The Wildcard Element

Definition 2.2.

Let the magnitude-wildcard space 𝕄∗=𝕄∪{∗}\mathbb{M}^{*}=\mathbb{M}\cup\{\ast\} be the set of magnitudes along with a wildcard element, ∗\ast.

The wildcard element corresponds to several different concepts, each appearing in a unique discipline:

  • •

    The NaN, or Not a Number33 3 “Not a Number” may have been an unfortunate naming choice because it actually represents any number! value in the IEEE standard for floating point arithmetic[10].

  • •

    The indeterminate form 00\frac{0}{0} in arithmetic.

  • •

    The wildcard pattern used in pattern matching and regular expressions in type theory and computer science

The following properties on ∗\ast allow multiplication of any two magnitude-wildcard values.

0⋅∞=∗∗⋅m=∗0\cdot\infty=\ast\qquad\ast\cdot m=\ast

We may also define ∗+m=∗\ast+m=\ast, but there is an argument to be made that because there are no negative numbers in this system, ∗+m\ast+m should be a new pattern that only matches magnitudes greater than or equal to mm. For our purposes, the question of how to handle this can be deferred.

2.3 The Matching Relation

Definition 2.3.

The matching relation :⁣≅:\cong is a binary relation on 𝕄∗\mathbb{M}^{*}. The statement m1:≅m2m_{1}:\cong m_{2} is read “m1m_{1} is matched by m2m_{2}” and is true when either m1m_{1} equals m2m_{2} or m2m_{2} is a wildcard.

m1:≅m2⟺(m1=m2)∨(m2=∗)m_{1}:\cong m_{2}\Longleftrightarrow(m_{1}=m_{2})\vee(m_{2}=\ast)

The left hand side of a matching relation is the parameter and the right hand side is the constraint. The wildcard element represents every single value, but it cannot be represented by any specific value. It also represents a loss of information about the parameter.

A few lemmas quickly follow.

Lemma 2.1.

If a magnitude matches a non-wildcard element, then the two values are equal.

m1:≅m2∧m2≠∗⟹m1=m2m_{1}:\cong m_{2}\wedge m_{2}\neq\ast\Longrightarrow m_{1}=m_{2}
Lemma 2.2.

Every element is matched by the wildcard element. m:≅∗m:\cong\ast

Lemma 2.3.

The wildcard element is matched only by itself. ∗:≅m⟹m=∗\ast:\cong m\Longrightarrow m=\ast

The matching relation looks a lot like equality and in many cases it is, but it doesn’t have all of the same properties.

Theorem 2.4.

The matching relation is reflexive, transitive, and anti-symmetric but not symmetric.

Proof.

Reflexive is obvious: m:≅m⟺(m=m)∨(m=∗)m:\cong m\Longleftrightarrow(m=m)\vee(m=\ast)

The transitive property states that for all m1,m2,m3m_{1},m_{2},m_{3} in 𝕄\mathbb{M}, if m1:≅m2m_{1}:\cong m_{2} and m2:≅m3m_{2}:\cong m_{3}, then m1:≅m3m_{1}:\cong m_{3}.

Assume that m1:≅m2m_{1}:\cong m_{2} and m2:≅m3m_{2}:\cong m_{3}. If none of these values are the wildcards, then by lemma 2.1, they are all equal and m1:≅m3m_{1}:\cong m_{3}. If m1=∗m_{1}=\ast then by lemma 2.3, m2=∗m_{2}=\ast and finally m3=∗m_{3}=\ast. In other words, if any of the three values are ∗\ast, then m3=∗m_{3}=\ast. By lemma 2.2, the theorem holds.

The anti-symmetric property states that if m1:≅m2m_{1}:\cong m_{2} and m2:≅m1m_{2}:\cong m_{1}, then m1=m2m_{1}=m_{2}.

If both m1m_{1} and m2m_{2} match each other, then (m1=m2)∨(m2=∗)(m_{1}=m_{2})\vee(m_{2}=\ast) and (m2=m1)∨(m1=∗)(m_{2}=m_{1})\vee(m_{1}=\ast). By boolean factorization, this means that (m1=m2)∨((m2=∗)∧(m1=∗))(m_{1}=m_{2})\vee((m_{2}=\ast)\wedge(m_{1}=\ast)). On both sides of the conjunction, m1=m2m_{1}=m_{2}.

The matching relation is not symmetric because m:≅∗m:\cong\ast but ∗:≇m\ast:\ncong m for any non-wildcard mm. ∎

3 Categorical Distributions

Let Ω\Omega be a set of mutually exclusive outcomes44 4 Each outcome could be thought of as a possible result of a random trial, or a possible value for an unknown variable. We assume that Ω\Omega is finite and there are KK outcomes or categories so that |Ω|=K|\Omega|=K.

Definition 3.1.

A categorical distribution on a Ω\Omega is a function P:Ω→[0,1]P:\Omega\rightarrow[0,1] such that ∑h∈ΩP⁡(h)=1\sum_{h\in\Omega}P(h)=1

The set of all categorical distributions of size KK can be embedded in ℝK\mathbb{R}^{K} as a (K-1)-dimensional object called a simplex (see figure 1). For example, if K=3K=3, the resulting space of categorical distributions is an equilateral triangle embedded in ℝ3\mathbb{R}^{3} connecting the points (1, 0, 0), (0, 1, 0), and (0, 0, 1).

P0P_{0}P1P_{1}(0,1)(1,0)
P0P_{0}P1P_{1}P2P_{2}(1,0,0)(0,0,1)(0,1,0)
(0,1,0,0)(0,0,0,1)(1,0,0,0)(0,0,1,0)
Figure 1: An illustration of the probability simplex for K = 2, 3, and 4. These objects are respectively, a segment embedded in ℝ2\mathbb{R}^{2}, an equilateral triangle embedded in ℝ3\mathbb{R}^{3}, and a normal tetrahedron embedded in ℝ4\mathbb{R}^{4}. We make no attempt to visualize the 4D space that contains the tetrahedron.

3.1 Events

An event is a set of outcomes, and by convention ℱ\mathcal{F} is the set of all possible events. In general, ℱ\mathcal{F} is not the entire power set of Ω\Omega, but when Ω\Omega is finite we can consider any subset e⊆Ωe\subseteq\Omega to be an event55 5 We need not concern ourselves with defining a σ\sigma-algebra of measurable sets..

Definition 3.1 creates a probability function that assigns values on individual outcomes. We now define the probability function on an event, which corresponds to the probability that any one of its outcomes occur. Looking at probability on the event level rather than the outcome level is a crucial insight in the development of probability theory (and measure theory more generally). Even though the process is far simpler for finite distributions, we must pay attention to this layer in order for the framework to generalize.

Definition 3.2.

For all ee in ℱ\mathcal{F},

P⁡(e)=∑h∈eP⁡(h).P(e)=\sum_{h\in e}{P(h)}.

PP acts on either outcomes or events using the convention P⁡({h})=P⁡(h)P(\{h\})=P(h). Ω\Omega is itself an event; the universal event of all outcomes, with probability 1.

P⁡(Ω)=∑h∈ΩP⁡(h)=1P(\Omega)=\sum_{h\in\Omega}{P(h)}=1

3.2 Relative Probability Function

A relative probability function, or RPF, measures the probability of one event with respect to another. For example, we may wish to talk about an event that is “twice as likely” as another, even if we don’t know the absolute probability of either event. In a sense, all probability is relative and conditional because all probabilistic statements come with underlying assumptions.

We continue to use P to represent the RPF but now with two inputs instead of one. The expression P⁡(e1,e2)P(e_{1},e_{2}) can be read as the probability of e1e_{1} relative to e2e_{2}.

P:ℱ×ℱ→𝕄∗P:\mathcal{F}\times\mathcal{F}\rightarrow\mathbb{M}^{*}

We define relative probability in terms of absolute probability as a ratio, in the style of Kolmogorov.

Definition 3.3.

The relative probability of events e1e_{1} and e2e_{2} on a categorical distribution PP is as follows:

P⁡(e1,e2)=P⁡(e1)P⁡(e2)P(e_{1},e_{2})=\frac{P(e_{1})}{P(e_{2})}

If P⁡(e1)=P⁡(e2)=0P(e_{1})=P(e_{2})=0, then P(e1,e2)=∗P(e_{1},e_{2})=\ast, representing the motivational problem of zero-probability events being incomparable.

With absolute probability, information is lost at the boundaries of the simplex where the probability of several outcomes might be assigned a value of zero. For example, if Ω={a,b,c}\Omega=\{a,b,c\} with P⁡(a)=1P(a)=1 and P⁡(b)=P⁡(c)=0P(b)=P(c)=0, we cannot compare the probabilities of bb and cc by ratio as we can in the rest of the simplex.

This poses an interesting problem for limits.

Example 3.1.

Consider the following categorical distribution function, with parameter ϵ>0\epsilon>0:

P⁡(a)=1−ϵP⁡(b)=23​ϵP⁡(c)=13​ϵP(a)=1-\epsilon\qquad P(b)=\frac{2}{3}\epsilon\qquad P(c)=\frac{1}{3}\epsilon

This is clearly an absolute probability, and its limit as ϵ\epsilon goes to zero should be P⁡(a)=1,P⁡(b)=P⁡(c)=0P(a)=1,P(b)=P(c)=0. The fact that b is twice as likely as c is lost!

4 The Relative Probability Approach

In section 3.2, the relative probability function was derived from the absolute probability function. Here in section 4, we invert this process by starting with the relative probability as the fundamental building block.

4.1 Fundamental Axioms

Definition 4.1.

Let Ω\Omega be the set of outcomes, and P:Ω×Ω→𝕄∗P:\Omega\times\Omega\rightarrow\mathbb{M}^{*} be a function acting on two outcomes to produce a magnitude-wildcard. PP is a relative probability function on the outcomes of Ω\Omega if it obeys the 3 fundamental axioms of relative probability:

  1. (i)

    The identity axiom: P⁡(h,h)=1P(h,h)=1

  2. (ii)

    The inverse axiom: P⁡(h1,h2)=P​(h2,h1)−1P(h_{1},h_{2})=P(h_{2},h_{1})^{-1}

  3. (iii)

    The composition axiom: P⁡(h1,h3):≅P⁡(h1,h2)⋅P⁡(h2,h3)P(h_{1},h_{3}):\cong P(h_{1},h_{2})\cdot P(h_{2},h_{3})

P⁡(h1,h2)P(h_{1},h_{2}) represents the probability of h1h_{1} relative to h2h_{2}. Outcomes h1h_{1} and h2h_{2} are comparable if P(h1,h2)≠∗P(h_{1},h_{2})\neq\ast.

Let us pause for a moment to discuss how these axioms were selected. The composition axiom is doing most of the work, and it succinctly encodes how relative probability works. If AA is twice as likely as BB, and BB is 3 times as likely as CC, then AA had better be 6 times as likely as CC. If not, these relative probability assignments would have no meaning; they would just be numerical assignments without rhyme or reason66 6 Many of our political and economic forecasts come in this form..

The composition axiom is enough to show that the identity axiom works most of the time. For any two outcomes h1h_{1} and h2h_{2} we get through composition P⁡(h1,h2):≅P⁡(h1,h1)⋅P⁡(h1,h2)P(h_{1},h_{2}):\cong P(h_{1},h_{1})\cdot P(h_{1},h_{2}). As long as P⁡(h1,h2)P(h_{1},h_{2}) isn’t 00, ∞\infty, or ∗\ast, then we would have to conclude P⁡(h1,h1)=1P(h_{1},h_{1})=1 so long as h1h_{1} is comparable to itself.

But that doesn’t get us all the way there! There are still scenarios where P(h,h)=∗P(h,h)=\ast. The question that must be asked is: should the the self-comparisons in an outcome space contain information where there is a choice of values between 11 and ∗\ast? Such information would not be relevant to relative probability. Therefore, P⁡(h,h)P(h,h) can only have a single value and it must be 1. Hence, the necessity of the identity axiom.

Composition and identity can actually be combined into a single axiom about composition paths. It’s a bit unwieldy, but nevertheless interesting.

Proposition 4.1 (Path Composition).

Given a non-empty list of NN outcomes h0,h1,h2,…,hN−1h_{0},h_{1},h_{2},...,h_{N-1},

P⁡(h0,hN−1):≅∏k=0N−2P⁡(hk,hk+1)P(h_{0},h_{N-1}):\cong\prod_{k=0}^{N-2}P(h_{k},h_{k+1})

In this case, P⁡(h0,h0)P(h_{0},h_{0}) would be matched by the empty product, which is 1.

The inverse axiom is nearly redundant as well. Since P⁡(h0,h0):≅P⁡(h0,h1)⋅P⁡(h1,h0)P(h_{0},h_{0}):\cong P(h_{0},h_{1})\cdot P(h_{1},h_{0}), the terms in the constraint look like they must be inverses. But without stating the axiom explicitly, there could be a case where P⁡(h0,h1)P(h_{0},h_{1}) is some non-wildcard magnitude like 2 but P⁡(h1,h0)P(h_{1},h_{0}) is ∗\ast. This shouldn’t be allowed because ∗\ast represents a lack of knowledge about a value, and we consider P⁡(h1,h0)P(h_{1},h_{0}) and P⁡(h1,h0)P(h_{1},h_{0}) to be the same piece of information but in reverse.

4.2 Examples

Now that the definition of relative probability is squared away, we can construct a library of examples for common RPFs that will serve as building blocks to describing common situations.

Definition 4.2.

The uniform RPF on any Ω\Omega considers each outcome equally likely. In other words, P⁡(h1,h2)=1P(h_{1},h_{2})=1 for every pair of outcomes.

Definition 4.3.

The indeterminate RPF has P(h1,h2)=∗P(h_{1},h_{2})=\ast for every pair of outcomes.

Definition 4.4.

A certain RPF contains a single outcome that has infinite probability relative to all other outcomes. Let hCh_{C} be the certain outcome with hC≠hh_{C}\neq h. Then P⁡(hC,h)=∞P(h_{C},h)=\infty. The relative probability of the other K−1K-1 outcomes could be anything.

Definition 4.5.

The empty RPF has no outcomes because K=0K=0, and therefore the function PP has no valid inputs.

It is surprising that there is still an RPF with Ω=∅\Omega=\varnothing. This is not the case for absolute distributions where such a function does not exist (because with no outcomes, they cannot sum to 1).

Definition 4.6.

The unit RPF has a single outcome where K=1K=1 and Ω=h\Omega=h. There is only one such RPF where P⁡(h,h)=1P(h,h)=1.

The unit RPF is both uniform and certain. This corresponds to the absolute case where the probability of the single outcome must be 1.

Definition 4.7.

Let PP be an RPF with K outcomes labeled (h0,h1,…,hK−1)(h_{0},h_{1},...,h_{K-1}). PP is a finite geometric RPF with ratio rr if the relative probabilities of each outcome with its neighbor is always rr. In other words, for all i∈(0,1,…,K−2)i\in(0,1,...,K-2),

P⁡(hi+1,hi)=rP(h_{i+1},h_{i})=r

When rr is 0 or ∞\infty, we can call this the limit finite geometric RPF.

Finally, to include an example that is both common and has powerful applications, the relative version of the binomial distribution can be defined as follows:

Definition 4.8.

A binomial distribution has a sample size nn, and a probability of success pp. The RPF has outcome space Ω={0,1,2,…,n}\Omega=\{0,1,2,...,n\} and thus K=n+1K=n+1. It is given as follows:

P⁡(h1,h2)=h2!​(n−h2)!h1!​(n−h1)!​(p1−p)h1−h2P(h_{1},h_{2})=\frac{h_{2}!(n-h_{2})!}{h_{1}!(n-h_{1})!}\left(\frac{p}{1-p}\right)^{h_{1}-h_{2}}

5 Concepts for Relative Probability Functions

We defined the relative probability function in section 4 with the fundamental axioms and have constructed some examples. Because new situations arise that do not occur in the Kolmogorov model, we also need to define some new vocabulary.

Figure 2 gives us a roadmap of these new concepts and their relationship to each other.

Relative Probability Function Definition 4.1Anchored RPFDefinition 5.3Empty RPFDefinition 4.5Indeterminate RPFDefinition 4.3Non-emptyTotally Comparable RPFDefinition 5.1Finite Limit Geometric Distribution,K>2K>2Definition 4.7Absolute ProbabilityFunctionDefinition 3.1UnionFacet ProbabilityRPF(single impossible event)Totally Mutually PossibleRPFDefinition 5.6(Non-Empty)UniformDefinition 4.2UnitDistributionDefinition 4.6Finite GeometricDistributionDefinition 4.7BinomialDistributionDefinition 4.8Limit FiniteGeometricwith K=2K=2Lemma 5.4Theorem 5.3Theorem 5.2Matched ByTheorem 5.5
Figure 2: This is a roadmap for all of the sub-types of relative probability functions and their relationship to one another.

5.1 Absolute Probability as a Special Case

Fortunately, the absolute probability function is a special case of an RPF (through definition 3.3), defined by P⁡(h1,h2)=P⁡(h1)P⁡(h2)P(h_{1},h_{2})=\frac{P(h_{1})}{P(h_{2})} with the exception that if P⁡(h)=0P(h)=0, then P⁡(h,h)=1P(h,h)=1 instead of ∗\ast in order to satisfy the identity axiom.

Theorem 5.1.

A relative probability function derived from an absolute probability function satisfies the fundamental axioms.

Proof.

For the identity axiom, we have P⁡(h,h)=1P(h,h)=1 built into the definition.

For the inverse axiom, we can see that for outcomes h1h_{1} and h2h_{2}, P⁡(h1,h2)=P⁡(h1)P⁡(h2)=(P⁡(h2)P⁡(h1))−1=P​(h2,h1)−1P(h_{1},h_{2})=\frac{P(h_{1})}{P(h_{2})}=\left(\frac{P(h_{2})}{P(h_{1})}\right)^{-1}=P(h_{2},h_{1})^{-1}.

For composition, we must show that for all outcomes h1,h2,h3h_{1},h_{2},h_{3}, P⁡(h1,h3):≅P⁡(h1,h2)⋅P⁡(h2,h3)P(h_{1},h_{3}):\cong P(h_{1},h_{2})\cdot P(h_{2},h_{3})

Start with the case that P⁡(h2)≠0P(h_{2})\neq 0. Then P⁡(h1,h2)⋅P⁡(h2,h3)=P⁡(h1)P⁡(h2)​P⁡(h2)P⁡(h3)=P⁡(h1)P⁡(h3)=P⁡(h1,h3)P(h_{1},h_{2})\cdot P(h_{2},h_{3})=\frac{P(h_{1})}{P(h_{2})}\frac{P(h_{2})}{P(h_{3})}=\frac{P(h_{1})}{P(h_{3})}=P(h_{1},h_{3}). When P⁡(h2)=0P(h_{2})=0, P(h1,h2)⋅P(h2,h3)=P⁡(h1)P⁡(h2)P⁡(h2)P⁡(h3)=∗P(h_{1},h_{2})\cdot P(h_{2},h_{3})=\frac{P(h_{1})}{P(h_{2})}\frac{P(h_{2})}{P(h_{3})}=\ast. Because ∗\ast matches everything, then the matching statement holds. Because it holds in both when P⁡(h2)=0P(h_{2})=0 and P⁡(h2)≠0P(h_{2})\neq 0, the theorem holds true always. ∎

5.2 Comparability, Possibility, and Anchors

Definition 5.1.

A relative probability function is totally comparable if every pair of outcomes are comparable.

Theorem 5.2.

An absolute probability function is totally comparable if and only if P⁡(h)=0P(h)=0 for at most one outcome.

Proof.

Let P be an absolute probability function with distinct outcomes h1h_{1} and h2h_{2}. If P⁡(h1)=P⁡(h2)=0P(h_{1})=P(h_{2})=0, then P(h1,h2)=00=∗P(h_{1},h_{2})=\frac{0}{0}=\ast. If only outcome h1h_{1} is assigned 0, then P⁡(h1,h1)=1P(h_{1},h_{1})=1, P⁡(h1,h2)=0P(h_{1},h_{2})=0, and P⁡(h2,h1)=∞P(h_{2},h_{1})=\infty. Any other pairing that does not involve h1h_{1} will be the quotient of two positive numbers, and thus not ∗\ast. ∎

Once it is established that outcomes are comparable to one another, they form groupings. If an outcomes h1h_{1} and h2h_{2} both have probability 0 (or ∞\infty) with respect to h3h_{3} then P⁡(h1,h2)P(h_{1},h_{2}) is not predetermined by composition. This leads us to the concept of possibility and impossibility.

Definition 5.2.

Outcome h1h_{1} is impossible with respect to h2h_{2} if P⁡(h1,h2)=0P(h_{1},h_{2})=0. Outcome h1h_{1} is possible with respect to h2h_{2} if it is comparable and P⁡(h1,h2)>0P(h_{1},h_{2})>0.

Anchor outcomes are those outcomes that have a non-zero absolute probability. The anchoring of a distribution ensures that it is well behaved.

Definition 5.3.

An anchor of an RPF is an outcome that is possible with respect to every other outcome. An RPF that has at least 1 anchor is called an anchored RPF.

Theorem 5.3.

All absolute probability distributions are anchored.

Proof.

Let P be an absolute probability distribution on Ω\Omega. Because ∑h∈ΩP⁡(h)=1\sum_{h\in\Omega}P(h)=1, there must be at least one hh such that P⁡(h)>0P(h)>0. Therefore, for any comparison outcome h′h^{\prime}, P⁡(h,h′)=P⁡(h)P⁡(h′)>0P(h,h^{\prime})=\frac{P(h)}{P(h^{\prime})}>0. ∎

Lemma 5.4.

Every non-empty, totally comparable RPF is anchored.

Proof.

Let PP be a non-empty and totally comparable RPF. To prove by contradiction, assume that arbitrary outcome hh is not an achor, and therefore there exists another outcome h′h^{\prime} such that P⁡(h,h′)=0P(h,h^{\prime})=0.

Create a function f:Ω→Ωf:\Omega\rightarrow\Omega that finds a potential h′=f⁡(h)h^{\prime}=f(h) so that P⁡(h,f⁡(h))=0P(h,f(h))=0.

Let fnf^{n} be the function ff applied n times. Then P​(h,fn​(h))=0P(h,f^{n}(h))=0 for all n greater than 0. This is by induction because the case of n=1n=1 was assumed above, and for the inductive step

P(h,fn+1(h)):≅P(h,fn(h))⋅P(fn(h),f(fn(h))=0⋅0=0P(h,f^{n+1}(h)):\cong P(h,f^{n}(h))\cdot P(f^{n}(h),f(f^{n}(h))=0\cdot 0=0

Because Ω\Omega is finite, repeated applications of ff on hh must eventually return to an outcome that has already been visited. In more rigorous terms, there exists an NN such that fN​(h)=fi​(h)f^{N}(h)=f^{i}(h) for some i<Ni<N.

This is a contradiction because P⁡(fi​(h),fN​(h))P(f^{i}(h),f^{N}(h)) should equal 0 by the argument above, but 1 by the identity axiom. ∎

A totally comparable RPF contains the maximum amount of information about the relative probability of its outcomes. Some RPFs have less information but are nevertheless consistent with those that have more. The following definition describes this relationship.

Definition 5.4.

Let P1P_{1} and P2P_{2} be relative probability functions. P1P_{1} is matched by P2P_{2} if and only if for all outcomes h1h_{1} and h2h_{2},

P1​(h1,h2):≅P2​(h1,h2).P_{1}(h_{1},h_{2}):\cong P_{2}(h_{1},h_{2}).
Theorem 5.5.

Every anchored RPF is matched by an absolute probability function, given by the following equation where aa is an anchor outcome.

P⁡(h)=P⁡(h,a)∑h′∈ΩP⁡(h′,a)P(h)=\frac{P(h,a)}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},a)}
Proof.

We need to show that P⁡(h)P(h) is a valid absolute probability function, and that it matches the original RPF.

Because aa is an anchor element, P⁡(h′,a)<∞P(h^{\prime},a)<\infty. Therefore the sum ∑h′∈ΩP⁡(h′,a)<∞\sum_{h^{\prime}\in\Omega}P(h^{\prime},a)<\infty. ∑h′∈ΩP⁡(h′,a)\sum_{h^{\prime}\in\Omega}P(h^{\prime},a) is also >0>0, because it includes the term P⁡(a,a)=1P(a,a)=1. The numerator P⁡(h,a)P(h,a) is also <∞<\infty. Therefore, P⁡(h)∉{∞,∗}P(h)\notin\{\infty,\ast\}.

We next check that the values of P⁡(h)P(h) sum to 1 as follows:

∑h∈ΩP⁡(h)=∑h∈ΩP⁡(h,a)∑h′∈ΩP⁡(h′,a)=∑h∈ΩP⁡(h,a)∑h′∈ΩP⁡(h′,a)=1\sum_{h\in\Omega}P(h)=\sum_{h\in\Omega}\frac{P(h,a)}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},a)}=\frac{\sum_{h\in\Omega}P(h,a)}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},a)}=1

Cancellation of these equal sums is justified because we have shown previously that they cannot be 00 or ∞\infty.

Therefore, P⁡(h)P(h) is a valid absolute probability function. We show that the RPF is matched by it though the following calculation:

P⁡(h1,h2):≅P⁡(h1,a)⋅P⁡(a,h2)=P⁡(h1,a)∑h′∈ΩP⁡(h′,a)÷P⁡(h2,a)∑h′∈ΩP⁡(h′,a)=P⁡(h1)P⁡(h2)P(h_{1},h_{2}):\cong P(h_{1},a)\cdot P(a,h_{2})=\frac{P(h_{1},a)}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},a)}\div\frac{P(h_{2},a)}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},a)}=\frac{P(h_{1})}{P(h_{2})} (1)

∎

5.3 Mutual Possibility

Definition 5.5.

Outcomes h1h_{1} and h2h_{2} are mutually possible if they are comparable and 0<P⁡(h1,h2)<∞0<P(h_{1},h_{2})<\infty. In other words, h1h_{1} and h2h_{2} are each possible with respect to the other.

Theorem 5.6.

Mutual possibility is an equivalence relation, being reflexive, symmetric and transitive.

Proof.

Reflexive: P⁡(h,h)=1P(h,h)=1 by the identity axiom.

Symmetric: P⁡(h1,h2)=P​(h2,h1)−1P(h_{1},h_{2})=P(h_{2},h_{1})^{-1}, which means that each can be in {0,∞,∗}\{0,\infty,\ast\} if and only if the other is as well.

Transitive: The composition axiom states that P⁡(h1,h3):≅P⁡(h1,h2)⋅P⁡(h2,h3)P(h_{1},h_{3}):\cong P(h_{1},h_{2})\cdot P(h_{2},h_{3}). If the factors in the constraint (P⁡(h1,h2)P(h_{1},h_{2}) and P⁡(h2,h3)P(h_{2},h_{3})) are positive and finite, then their product is also positive and finite. ∎

Definition 5.6.

An RPF is totally mutually possible if all of its outcomes77 7 This is one of the few definitions that cannot later be upgraded from outcomes to events. The empty event e={}e=\{\} for example will be impossible with respect to any outcome by theorem 6.2. are mutually possible or equivalently are all anchors.

It is helpful to make diagrams of possibility and impossibility through a directed graph. In these graphs, each outcome is represented by a point, and an arrow from A to B means that B is possible with respect to A. A bidirectional arrow means that A and B are mutually possible. Totally mutually possible RPFs have a simple diagram where all the outcomes are connected as in figure 3.

Figure 3: A totally mutually possible RFP has - unsurprisingly - a complete graph of mutual possibility.
Theorem 5.7.

A non-empty totally mutually possible RPF is equal to an absolute probability function.

Proof.

If PP is totally mutually possible, then all of its outcomes are anchors. Therefore, we can use theorem 5.5 to find a matching absolute probability function

P⁡(h)=P⁡(h,a)∑h′∈ΩP⁡(h′,a)P(h)=\frac{P(h,a)}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},a)}

Because every element of Ω\Omega is an anchor, we can let a=ha=h and get

P⁡(h)=P⁡(h,h)∑h′∈ΩP⁡(h′,h)=1∑h′∈ΩP⁡(h′,h)P(h)=\frac{P(h,h)}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},h)}=\frac{1}{\sum_{h^{\prime}\in\Omega}P(h^{\prime},h)}

Theorem 5.5 states that P⁡(h1,h2):≅P⁡(h1)P⁡(h2)P(h_{1},h_{2}):\cong\frac{P(h_{1})}{P(h_{2})}, but since the constraint P⁡(h1)P⁡(h2)\frac{P(h_{1})}{P(h_{2})} is never ∗\ast, they must be equal. ∎

5.4 Possibility Classes

In order to analyze the general case of RPFs, we need to consider classes of mutual possibility.

Theorem 5.8.

The relationship of being possible is both reflexive and transitive, or in other words a preorder.

Proof.

It is reflexive because P⁡(h,h)=1P(h,h)=1. If P⁡(h1,h2)>0P(h_{1},h_{2})>0 and P⁡(h2,h3)>0P(h_{2},h_{3})>0 then their product is also greater than zero, and by composition, equal to P⁡(h1,h3)P(h_{1},h_{3}). Thus h1h_{1} is also possible with respect to h3h_{3}. ∎

If we consider the equivalence classes of mutual possibility and their relationship to one another, then we have a partial order. Figures 4 and 5 are both examples of a graph of outcomes grouped by mutually possible equivalence classes. Figure 4 is anchored while figure 5 is not.

Figure 4: A diagram of an anchored RPF with its mutually possible classes. Each point is an outcome, and each circle is mutually possible class. The anchor class (shaded) is the maximal class in the partial order.
Figure 5: A diagram for a single RPF that is not anchored. We cannot turn this into an absolute probability function.
Theorem 5.9.

Consider two distinct equivalence classes of mutually possible outcomes in an RPF and call them C1,C2⊆ΩC_{1},C_{2}\subseteq\Omega. Any pair of outcomes h1∈C1h_{1}\in C_{1} and h2∈C2h_{2}\in C_{2} will have the same relative probability P⁡(h1,h2)P(h_{1},h_{2}) as any other pair of outcomes in C1×C2C_{1}\times C_{2}.

Proof.

P⁡(h1,h2)P(h_{1},h_{2}) must be in {0,∞,∗}\{0,\infty,\ast\} because otherwise h1h_{1} and h2h_{2} would be mutually possible and in the same equivalence class. Let h1′∈C1h^{\prime}_{1}\in C_{1} and h2′∈C2h^{\prime}_{2}\in C_{2} be alternative representatives from equivalences classes C1C_{1} and C2C_{2}. Then 0<P⁡(h1′,h1)<∞0<P(h^{\prime}_{1},h_{1})<\infty and 0<P⁡(h2,h2′)<∞0<P(h_{2},h^{\prime}_{2})<\infty due to the definition of mutual comparability. Thus with composition we get

P⁡(h1′,h2′):≅P⁡(h1′,h1)⋅P⁡(h1,h2)⋅P⁡(h2,h2′)=P⁡(h1,h2)P(h^{\prime}_{1},h^{\prime}_{2}):\cong P(h^{\prime}_{1},h_{1})\cdot P(h_{1},h_{2})\cdot P(h_{2},h^{\prime}_{2})=P(h_{1},h_{2})

∎

Finally, we look at totally comparable RPFs, where mutually possible components form a total order (see figure 6). We’ve established in the proof of theorem 5.9 that the relative probability of outcomes from one equivalence class to another are constant and either 0, ∞\infty or ∗\ast. If the RPF is totally comparable, then it can only be 0 or ∞\infty, providing a transitive, binary relation and thus a total order.

Figure 6: A diagram of a totally comparable RPF that is not totally mutually possible. The mutually possible components form a total order, with the anchored component on top and the graph forming a line.

6 From Outcomes to Events

Our next task is to upgrade PP to operate on the event level. This is more difficult than it seems. For example, we may wish to declare that the probability of event e1e_{1} with respect to e2e_{2} is going to be additive on e1e_{1} as follows:

P⁡(e1,e2)=∑h1∈e1P⁡(h1,e2)P(e_{1},e_{2})=\sum_{h_{1}\in e_{1}}P(h_{1},e_{2}) (2)

Equation 2 looks uncontroversial, but it actually contradicts the fundamental axioms! If we let e1=∅e_{1}=\varnothing, then we have an empty sum on the right hand side of the equation, and we get P⁡(∅,e2)=0P(\varnothing,e_{2})=0. Likewise, if we allow e2e_{2} to be empty, we get P⁡(e1,∅)=P​(∅,e1)−1=0−1=∞P(e_{1},\varnothing)=P(\varnothing,e_{1})^{-1}=0^{-1}=\infty. Both of these statements make sense until they collide in P⁡(∅,∅)=0=∞P(\varnothing,\varnothing)=0=\infty, and what’s worse is that this is also equal 1 under the identity axiom!

Another problem arises when an event is internally incomparable, meaning that it contains outcomes that are incomparable with each other. Perhaps there are interesting things we can say about such events, but here we will constrain ourselves to totally comparable RPFs in order to avoid such questions.

Definition 6.1.

Let PP be a totally comparable RPF. PP measures the probability of two events relative to each other using the following rules:

  1. (i)

    P⁡(e1,e2)P(e_{1},e_{2}) obeys the fundamental axioms of relative probability.

  2. (ii)

    P⁡(e1,e2)P(e_{1},e_{2}) sums over any reference outcome rr as follows:

    P⁡(e1,e2):≅∑h1∈e1P⁡(h1,r)∑h2∈e2P⁡(h2,r).P(e_{1},e_{2}):\cong\frac{\sum_{h_{1}\in e_{1}}P(h_{1},r)}{\sum_{h_{2}\in e_{2}}P(h_{2},r)}. (3)

Because we no longer have access to absolute probability, the best we can do is measure it relative to a reference outcome rr. This ratio might be indeterminate, so we use the matching relation instead of equality. Fortunately, we can show that there exists at least one reference outcome that will constrain P⁡(e1,e2)P(e_{1},e_{2}) in statement 3 so long as e1e_{1} and e2e_{2} are not both empty.

Proof.

Lemma 5.4 states that all totally comparable RPFs have anchor outcomes, and therefore (by the same argument) every event must contain internal anchors, defined as an anchor for the RPF that is restricted to outcomes in that event. Choose an internal anchor aa from one of the events, say e1e_{1}. Then the sum ∑h1∈e1P⁡(h1,a)\sum_{h_{1}\in e_{1}}P(h_{1},a) will be non-infinite88 8 For anchor a, P⁡(a,h)>0P(a,h)>0 but this also means that P⁡(h,a)<∞P(h,a)<\infty., and non-zero because P⁡(a,a)=1P(a,a)=1 is a term in the sum. Therefore, the constraint as a whole cannot be indeterminate. ∎

If both events are empty then they cannot have internal anchors, but by the identity axiom P⁡(∅,∅)=1P(\varnothing,\varnothing)=1.

These requirements again seem reasonable, but how can we know for sure that they provide a complete and consistent definition of P:ℱ×ℱ→𝕄P:\mathcal{F}\times\mathcal{F}\rightarrow\mathbb{M}? The following must be shown:

  1. (i)

    If two distinct reference outcomes for rr in statement 3 yield constraints on PP, then they must be equal.

  2. (ii)

    The constraint in statement 3 does not violate the fundamental axioms.

Proof.

For i:

Let r1r_{1} and r2r_{2} be distinct reference outcomes that constrain P⁡(e1,e2)P(e_{1},e_{2}). Then we want to check that

∑h1∈e1P⁡(h1,r1)∑h2∈e2P⁡(h2,r1)=∑h1∈e1P⁡(h1,r2)∑h2∈e2P⁡(h2,r2)\frac{\sum_{h_{1}\in e_{1}}P(h_{1},r_{1})}{\sum_{h_{2}\in e_{2}}P(h_{2},r_{1})}=\frac{\sum_{h_{1}\in e_{1}}P(h_{1},r_{2})}{\sum_{h_{2}\in e_{2}}P(h_{2},r_{2})} (4)

Neither expression is a wildcard, and none of the individual terms are either. The key to this argument is in the value of P⁡(r1,r2)P(r_{1},r_{2}).

Assume P⁡(r1,r2)=0P(r_{1},r_{2})=0.

If ∑h1∈e1P⁡(h1,r1)\sum_{h_{1}\in e_{1}}P(h_{1},r_{1}) is finite, then ∑h1∈e1P⁡(h1,r2)\sum_{h_{1}\in e_{1}}P(h_{1},r_{2}) must be matched by ∑h1∈e1P⁡(h1,r1)⋅P⁡(r1,r2)\sum_{h_{1}\in e_{1}}P(h_{1},r_{1})\cdot P(r_{1},r_{2}) which is 0. By the same argument, ∑h2∈e2P⁡(h2,r2)\sum_{h_{2}\in e_{2}}P(h_{2},r_{2}) is also 0. Since they can’t both be zero, one of the sums involving r1r_{1} on the right hand side is infinite. Therefore P⁡(e1,e2)P(e_{1},e_{2}) is either ∞\infty or 0. Let’s say it is P⁡(e1,e2)=0P(e_{1},e_{2})=0. Then ∑h2∈e2P⁡(h2,r1)=∞\sum_{h_{2}\in e_{2}}P(h_{2},r_{1})=\infty and by the argument above ∑h1∈e1P⁡(h1,r2)=0\sum_{h_{1}\in e_{1}}P(h_{1},r_{2})=0. Because the right hand side is not ∗\ast - it must resolve to zero as well.

By analogous arguments, equation 4 must also hold when P⁡(e1,e2)=∞P(e_{1},e_{2})=\infty or P⁡(r1,r2)=∞P(r_{1},r_{2})=\infty or both.

Next assume that P⁡(r1,r2)∉{0,∞}P(r_{1},r_{2})\notin\{0,\infty\}. Multiply the left hand side of equation 4 by 1=P⁡(r1,r2)P⁡(r1,r2)1=\frac{P(r_{1},r_{2})}{P(r_{1},r_{2})} and distribute to get:

∑h1∈e1P⁡(h1,r1)⋅P⁡(r1,r2)∑h2∈e2P⁡(h2,r1)⋅P⁡(r1,r2)=∑h1∈e1P⁡(h1,r2)∑h2∈e2P⁡(h2,r2)\frac{\sum_{h_{1}\in e_{1}}P(h_{1},r_{1})\cdot P(r_{1},r_{2})}{\sum_{h_{2}\in e_{2}}P(h_{2},r_{1})\cdot P(r_{1},r_{2})}=\frac{\sum_{h_{1}\in e_{1}}P(h_{1},r_{2})}{\sum_{h_{2}\in e_{2}}P(h_{2},r_{2})}

For ii:

The identity, inverse, and composition axioms follow from the fact that statement 3 is a ratio with identical expressions for e1e_{1} in the numerator and e2e_{2} in the denominator. Therefore, if it resolves it is just a ratio of positive numbers, which follow the fundamental axioms. ∎

Theorem 6.1.

If events e1e_{1} and e2e_{2} are not both empty, the following formula calculates the relative probability of events:

P⁡(e1,e2)=∑h1∈e11∑h2∈e2P⁡(h2,h1).P(e_{1},e_{2})=\sum_{h_{1}\in e_{1}}\frac{1}{\sum_{h_{2}\in e_{2}}P(h_{2},h_{1})}.
Proof.

Find a suitable reference outcome rr and multiply by 1=P⁡(h1,r)P⁡(h1,r)1=\frac{P(h_{1},r)}{P(h_{1},r)}.

∑h1∈e11∑h2∈e2p⁡(h2,h1):≅∑h1∈e1P⁡(h1,r)∑h2∈e2P⁡(h2,h1)​P​(h1,r)=∑h1∈e1P⁡(h1,r)∑h2∈e2P⁡(h2,r)\sum_{h_{1}\in e_{1}}\frac{1}{\sum_{h_{2}\in e_{2}}p(h_{2},h_{1})}:\cong\sum_{h_{1}\in e_{1}}\frac{P(h_{1},r)}{\sum_{h_{2}\in e_{2}}P(h_{2},h_{1})P(h_{1},r)}=\frac{\sum_{h_{1}\in e_{1}}P(h_{1},r)}{\sum_{h_{2}\in e_{2}}P(h_{2},r)}

Since both P⁡(e1,e2)P(e_{1},e_{2}) and ∑h1∈e11∑h2∈e2p⁡(h2,h1)\sum_{h_{1}\in e_{1}}\frac{1}{\sum_{h_{2}\in e_{2}}p(h_{2},h_{1})} match ∑h1∈e1P⁡(h1,r)∑h2∈e2P⁡(h2,r)\frac{\sum_{h_{1}\in e_{1}}P(h_{1},r)}{\sum_{h_{2}\in e_{2}}P(h_{2},r)} which is not ∗\ast for appropriate reference r, they must be equal. ∎

We then derive the absolute probability function as

P⁡(e)=P⁡(e,Ω)=∑h∈e1∑h′∈Ωp⁡(h′,h)P(e)=P(e,\Omega)=\sum_{h\in e}\frac{1}{\sum_{h^{\prime}\in\Omega}p(h^{\prime},h)}
Theorem 6.2.

The empty event ∅\varnothing has probability 0 relative to any non-empty event.

Proof.

Let ee be a non-empty event, and let hh be an outcome in ee.

P⁡(∅,e):≅∑h1∈∅P⁡(h1,h)∑h2∈eP⁡(h2,h)=0∑h2∈eP⁡(h2,h)P(\varnothing,e):\cong\frac{\sum_{h_{1}\in\varnothing}P(h_{1},h)}{\sum_{h_{2}\in e}P(h_{2},h)}=\frac{0}{\sum_{h_{2}\in e}P(h_{2},h)}

The sum ∑h2∈eP⁡(h2,h)\sum_{h_{2}\in e}P(h_{2},h) cannot itself be zero because P⁡(h,h)P(h,h) is one of its terms. Therefore, P⁡(∅,e)=0P(\varnothing,e)=0 ∎

7 Composing Relative Probability Functions

Let P0,P1,…,PK−1P_{0},P_{1},...,P_{K-1} be relative probability functions on outcome spaces Ω0,Ω1,…,ΩK−1\Omega_{0},\Omega_{1},...,\Omega_{K-1} respectively so that each PkP_{k} is a function of type Ωk×Ωk→𝕄∗\Omega_{k}\times\Omega_{k}\rightarrow\mathbb{M}^{\ast}.

We can combine all of these relative probability functions together with a top level probability function P⊤P_{\top} (pronounced “P-Top”) with outcome space Ω⊤={Ω0,Ω1,…​ΩK−1}\Omega_{\top}=\{\Omega_{0},\Omega_{1},...\Omega_{K-1}\}. The outcome space is hierarchical as shown in figure 7.

Ω⊤\Omega_{\top}Ω0\Omega_{0}h0,0h_{0,0}h0,1h_{0,1}…h0,|Ω0|−1h_{0,|\Omega_{0}|-1}Ω1\Omega_{1}h1,0h_{1,0}…………ΩK−1\Omega_{K-1}…hK−1,|ΩK−1|−1h_{K-1,|\Omega_{K-1}|-1}
Figure 7: A tree diagram for a set of RPFs being composed by a top-level RPF.

Now let Ω\Omega be the set of all outcomes Ω0∪Ω1∪…​ΩK−1\Omega_{0}\cup\Omega_{1}\cup\dots\Omega_{K-1}. We can create a new relative probability function PP acting on outcome space Ω\Omega with the following rules:

  • •

    If the two outcomes fall under the same component, then their relative probabilities do not change:

    P⁡(hk,i,hk,j)=Pk​(hk,i,hk,j)P(h_{k,i},h_{k,j})=P_{k}(h_{k,i},h_{k,j}) (5)
  • •

    If the two outcomes fall under different components, then their relative probabilities are given as follows.

    P⁡(hk1,i,hk2,j)=Pk1​(hk1,i,Ωk1)⋅P⊤​(Ωk1,Ωk2)⋅Pk2​(Ωk2,hk2,j)P(h_{k_{1},i},h_{k_{2},j})=P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}})\cdot P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})\cdot P_{k_{2}}(\Omega_{k_{2}},h_{k_{2},j}) (6)

Note the use of composition to traverse up and down the tree. One could of course imagine this tree being many levels, and having a different height for each branch.

Theorem 7.1.

Composed function PP as defined above respects the fundamental axioms.

Proof.

The axioms are already obeyed by each component internally, but they also need to hold when the two inputs to PP are in different components. An outcome is always in the same component as itself, so the identity axiom immediately follows from equation 5.

The inverse law can be proven by calculation.

P​(hk1,i,hk2,j)−1\displaystyle P(h_{k_{1},i},h_{k_{2},j})^{-1} =(Pk1​(hk1,i,Ωk1)⋅P⊤​(Ωk1,Ωk2)⋅Pk2​(Ωk2,hk2,j))−1\displaystyle=(P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}})\cdot P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})\cdot P_{k_{2}}(\Omega_{k_{2}},h_{k_{2},j}))^{-1} (7)
=Pk1​(hk1,i,Ωk1)−1⋅P⊤​(Ωk1,Ωk2)−1⋅Pk2​(Ωk2,hk2,j)−1\displaystyle=P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}})^{-1}\cdot P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})^{-1}\cdot P_{k_{2}}(\Omega_{k_{2}},h_{k_{2},j})^{-1}
=Pk1​(Ωk1,hk1,i)⋅P⊤​(Ωk2,Ωk1)⋅Pk2​(hk2,j,Ωk2)\displaystyle=P_{k_{1}}(\Omega_{k_{1}},h_{k_{1},i})\cdot P_{\top}(\Omega_{k_{2}},\Omega_{k_{1}})\cdot P_{k_{2}}(h_{k_{2},j},\Omega_{k_{2}})
=Pk2​(hk2,j,Ωk2)⋅P⊤​(Ωk2,Ωk1)⋅Pk1​(Ωk1,hk1,i)\displaystyle=P_{k_{2}}(h_{k_{2},j},\Omega_{k_{2}})\cdot P_{\top}(\Omega_{k_{2}},\Omega_{k_{1}})\cdot P_{k_{1}}(\Omega_{k_{1}},h_{k_{1},i})
=P⁡(hk2,j,hk1,i)\displaystyle=P(h_{k_{2},j},h_{k_{1},i})

Composition can be shown similarly - now naming the 3 separate indices in components k1,k2,k3k_{1},k_{2},k_{3} as i1,i2,i3i_{1},i_{2},i_{3} respectively.

P⁡(hk1,i1,hk2,i2)⋅P⁡(hk2,i2,hk3,i3)\displaystyle P(h_{k_{1},i_{1}},h_{k_{2},i_{2}})\cdot P(h_{k_{2},i_{2}},h_{k_{3},i_{3}}) (8)
=Pk1​(hk1,i1,Ωk1)⋅P⊤​(Ωk1,Ωk2)⋅Pk2​(Ωk2,hk2,i2)⋅Pk2​(hk2,i2,Ωk2)⋅P⊤​(Ωk2,Ωk3)⋅Pk3​(Ωk3,hk3,i3)\displaystyle=P_{k_{1}}(h_{k_{1},i_{1}},\Omega_{k_{1}})\cdot P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})\cdot{\color[rgb]{1,0,0}P_{k_{2}}(\Omega_{k_{2}},h_{k_{2},i_{2}})\cdot P_{k_{2}}(h_{k_{2},i_{2}},\Omega_{k_{2}})}\cdot P_{\top}(\Omega_{k_{2}},\Omega_{k_{3}})\cdot P_{k_{3}}(\Omega_{k_{3}},h_{k_{3},i_{3}})
:≅Pk1​(hk1,i1,Ωk1)⋅P⊤​(Ωk1,Ωk2)⋅P⊤​(Ωk2,Ωk3)⋅Pk3​(Ωk3,hk3,i3)\displaystyle:\cong P_{k_{1}}(h_{k_{1},i_{1}},\Omega_{k_{1}})\cdot{\color[rgb]{1,0,0}P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})\cdot P_{\top}(\Omega_{k_{2}},\Omega_{k_{3}})}\cdot P_{k_{3}}(\Omega_{k_{3}},h_{k_{3},i_{3}})
:≅Pk1​(hk1,i1,Ωk1)⋅P⊤​(Ωk1,Ωk3)⋅Pk3​(Ωk3,hk3,i3)\displaystyle:\cong P_{k_{1}}(h_{k_{1},i_{1}},\Omega_{k_{1}})\cdot P_{\top}(\Omega_{k_{1}},\Omega_{k_{3}})\cdot P_{k_{3}}(\Omega_{k_{3}},h_{k_{3},i_{3}})
:≅Pk1​(hk1,i1,hk3,i3)\displaystyle:\cong P_{k_{1}}(h_{k_{1},i_{1}},h_{k_{3},i_{3}})

∎

Theorem 7.2.

PP is totally comparable if and only if the following are true:

  1. (i)

    P⊤P_{\top} and PkP_{k} for all k∈{0,1,…,K−1}k\in\{0,1,...,K-1\} are totally comparable.

  2. (ii)

    All components except at most one are totally mutually possible.

  3. (iii)

    If there is a component that is not totally mutually possible, then every element of P⊤P_{\top} is possible with respect to that component.

Proof.

First assume the conditions to show that the total comparability of PP follows.

If all the components are totally comparable, then any two outcomes in the same component are always going to be comparable in the overall RPF. We only need to prove that outcomes in different components are comparable. Starting with equation 6,

P⁡(hk1,i,hk2,j)=Pk1​(hk1,i,Ωk1)⋅P⊤​(Ωk1,Ωk2)⋅Pk2​(Ωk2,hk2,j)P(h_{k_{1},i},h_{k_{2},j})=P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}})\cdot P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})\cdot P_{k_{2}}(\Omega_{k_{2}},h_{k_{2},j}) (9)

The only way that P(hk1,i,hk2,j)=∗P(h_{k_{1},i},h_{k_{2},j})=\ast is if both 00 and ∞\infty are factors of the constraint Pk1​(hk1,i,Ωk1)⋅P⊤​(Ωk1,Ωk2)⋅Pk2​(Ωk2,hk2,j)P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}})\cdot P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})\cdot P_{k_{2}}(\Omega_{k_{2}},h_{k_{2},j})

Because there is at most one component that is not totally mutually possible by ii, we can say that either Pk1​(hk1,i,Ωk1)=0P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}})=0 or Pk2​(hk2,j,Ωk2)=0P_{k_{2}}(h_{k_{2},j},\Omega_{k_{2}})=0, or possibly neither, but not both.

Using equation 3 with the factor Pk1​(hk1,i,Ωk1)P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}}) and k1k_{1} itself as the reference outcome, we get

Pk1​(hk1,i,Ωk1):≅∑h1∈{k1}P⁡(h1,k1)∑h∈Ωk1P⁡(h2,k1)=1∑h∈Ωk1P⁡(h2,k1).P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}}):\cong\frac{\sum_{h_{1}\in\{k_{1}\}}P(h_{1},k_{1})}{\sum_{h\in\Omega_{k_{1}}}P(h_{2},k_{1})}=\frac{1}{\sum_{h\in\Omega_{k_{1}}}P(h_{2},k_{1})}.

The sum in the denominator cannot be zero since P⁡(k1,k1)=1P(k_{1},k_{1})=1 will be one of its terms. Therefore the factors Pk1​(hk1,i,Ωk1)P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}}) and Pk2​(hk2,j,Ωk2)P_{k_{2}}(h_{k_{2},j},\Omega_{k_{2}}) also must be finite.

If the term Pk1​(hk1,i,Ωk1)=0P_{k_{1}}(h_{k_{1},i},\Omega_{k_{1}})=0, then the only way the entire right hand side can be ∗\ast is if P⊤​(Ωk1,Ωk2)=∞P_{\top}(\Omega_{k_{1}},\Omega_{k_{2}})=\infty. But this can’t be true because by iii we assumed that Ωk2\Omega_{k_{2}} is possible with respect to Ωk1\Omega_{k_{1}}, the sole component with impossible outcomes. An analogous argument can be made if Pk2​(hk2,j,Ωk2)=0P_{k_{2}}(h_{k_{2},j},\Omega_{k_{2}})=0.

Therefore, the right hand side of the equation is not ∗\ast and PP is totally comparable.

In the opposite direction, we show that if any of the conditions are broken, then PP is not totally comparable. Breaking condition i would introduce an explicit ∗\ast into equation 6. If there are multiple components with impossible outcomes breaking condition ii, then it would introduce a 00 into the first term of equation 6 and an ∞\infty into the third term, yielding ∗\ast.

And finally, if only condition iii is broken, it would (with appropriate choices of h1h_{1} and h2h_{2}) introduce a 0 into the first term of equation 6 and an ∞\infty into the second term of equation 6.

Therefore, if any of these conditions are broken, PP is not totally comparable. ∎

8 Bayesian Inference on Relative Distributions

Bayesian inference is the practice of updating ones beliefs about the state of the world given new data.

A relative probability function represents a belief over the set of potential hypotheses in Ω\Omega, meaning that Bayesian inference can be performed on RPFs. The initial belief is called the prior and the updated belief is called the posterior. Once a formula for the posterior is is worked out, the practitioner might then want to search the hypothesis space Ω\Omega for an outcome that is either most likely (maximum a posteriori) or very good with respect to the posterior distribution. They might also wish to randomly sample one or more outcomes h∈Ωh\in\Omega from the posterior weighted according to the relative probabilities.

Almost all of these sampling and search methods rely on an algorithm that starts at one or more initial hypotheses and iterately updates using the relative probability of a current outcome and its nearby outcomes. Examples include hill climbing, simulated annealing, Newton-Raphson, Markov Chain Monte Carlo, and the No U-Turn sampler. The role of these algorithms in supervised machine learning has been previously discussed by Local Maximum Labs in “Sampling Bias Correction for Supervised Machine Learning[11]”. Because relative probability simplifies Bayes rule and is ideal for many sampling and selection algorithms, statisticians and engineers should consider using the RPF framework for these purposes.

Start with the Bayesian inference formula for conditional probability for h∈Ωh\in\Omega assuming that we receive data DD.

P⁡(h|D)=P⁡(D|h)⋅P⁡(h)P⁡(D)P⁡(D)=∑h∈ΩP⁡(D|h)⋅P⁡(h)P(h|D)=\frac{P(D|h)\cdot P(h)}{P(D)}\qquad P(D)=\sum_{h\in\Omega}P(D|h)\cdot P(h)

Now we convert to relative probability by looking at two hypotheses h1h_{1} and h2h_{2} and the ratio of their posterior probabilities.

P⁡(h1|D)P⁡(h2|D)=P⁡(D|h1)⋅P⁡(h1)P⁡(D)÷P⁡(D|h2)⋅P⁡(h2)P⁡(D)=P⁡(D|h1)⋅P⁡(h1)P⁡(D|h2)⋅P⁡(h2)\frac{P(h_{1}|D)}{P(h_{2}|D)}=\frac{P(D|h_{1})\cdot P(h_{1})}{P(D)}\div\frac{P(D|h_{2})\cdot P(h_{2})}{P(D)}=\frac{P(D|h_{1})\cdot P(h_{1})}{P(D|h_{2})\cdot P(h_{2})}

Notice that each component is now represented by a ratio. By making the appropriate substitutions, we can express this entirely in terms of RPFs.

For the ratio of prior probabilities, substitute the relative prior: P⁡(h1)P⁡(h2)→P⁡(h1,h2)\frac{P(h_{1})}{P(h_{2})}\rightarrow P(h_{1},h_{2})

For the ratio of posterior probabilities, substitute the relative posterior: P⁡(h1|D)P⁡(h2|D)→P⁡(h1,h2|D)\frac{P(h_{1}|D)}{P(h_{2}|D)}\rightarrow P(h_{1},h_{2}|D)

The liklihood ratio, also known as the bayes factor, may not look like a relative probability, but the formula to expand conditional probability suggests that it is:

P⁡(D|h1)P⁡(D|h2)=P⁡(D∩h1)P⁡(D)P⁡(D∩h2)P⁡(D)=P⁡(D∩h1)P⁡(D∩h2)\frac{P(D|h_{1})}{P(D|h_{2})}=\frac{\frac{P(D\cap h_{1})}{P(D)}}{\frac{P(D\cap h_{2})}{P(D)}}=\frac{P(D\cap h_{1})}{P(D\cap h_{2})}

Let PDP_{D} represent the likelihood ratio of the different hypotheses. The likelihood ratio PD​(h1,h2)P_{D}(h_{1},h_{2}) encodes a description of how the different hypotheses rate the likelihood of data.

The substitution for the likelihood ratio is now as follows: P⁡(D|h1)P⁡(D|h2)→PD​(h1,h2)\frac{P(D|h_{1})}{P(D|h_{2})}\rightarrow P_{D}(h_{1},h_{2})

These substitutions create a Bayes rule for relative probability:

P⁡(h1,h2|D)=PD​(h1,h2)​P​(h1,h2)P(h_{1},h_{2}|D)=P_{D}(h_{1},h_{2})P(h_{1},h_{2}) (10)

Bayesian inference is now reduced to a term-by-term multiplication of two different RPFs: PD​(h1,h2)P_{D}(h_{1},h_{2}) and P⁡(h1,h2)P(h_{1},h_{2}). Fortunately, the product of two RPFs also obey the fundamental axioms.

Theorem 8.1.

Let P1P_{1} and P2P_{2} be relative probability functions on Ω\Omega. Define P⁡(h1,h2)=P1​(h1,h2)⋅P2​(h1,h2)P(h_{1},h_{2})=P_{1}(h_{1},h_{2})\cdot P_{2}(h_{1},h_{2}). Then, PP is also an RPF because it obeys the fundamental axioms.

Proof.

Identity: P⁡(h1,h1)=P1​(h1,h1)​P2​(h1,h1)=1⋅1=1P(h_{1},h_{1})=P_{1}(h_{1},h_{1})P_{2}(h_{1},h_{1})=1\cdot 1=1

Inverse:

P⁡(h1,h2)=P1​(h1,h2)⋅P2​(h1,h2)=P1​(h2,h1)−1⋅P2​(h2,h1)−1=(P1​(h2,h1)⋅P2​(h2,h1))−1=P​(h2,h1)−1P(h_{1},h_{2})=P_{1}(h_{1},h_{2})\cdot P_{2}(h_{1},h_{2})=P_{1}(h_{2},h_{1})^{-1}\cdot P_{2}(h_{2},h_{1})^{-1}=(P_{1}(h_{2},h_{1})\cdot P_{2}(h_{2},h_{1}))^{-1}=P(h_{2},h_{1})^{-1}

Composition:

P⁡(h1,h2)​P​(h2,h3)=P1​(h1,h2)​P2​(h1,h2)​P1​(h2,h3)​P2​(h2,h3):≅P1​(h1,h3)​P2​(h1,h3)=P⁡(h1,h3)P(h_{1},h_{2})P(h_{2},h_{3})=P_{1}(h_{1},h_{2})P_{2}(h_{1},h_{2})P_{1}(h_{2},h_{3})P_{2}(h_{2},h_{3}):\cong P_{1}(h_{1},h_{3})P_{2}(h_{1},h_{3})=P(h_{1},h_{3})

∎

8.1 Degeneration of Mutually Possible Beliefs

The Cromwell rule in Bayesian inference cautions against setting the probability of outcomes equal to 0 in a prior belief. For relative probability, this corresponds to a preference for totally mutually possible RPFs. If an outcome were considered impossible with respect to another, this would be a permanent belief that cannot be changed through evidence.

Theorem 8.2.

If an outcome is impossible with respect to another outcome in the posterior distribution, it will either remain impossible or become incomparable in the posterior. In other words, if P⁡(h1,h2)=0P(h_{1},h_{2})=0, then P⁡(h1,h2|D)∈{0,∗}P(h_{1},h_{2}|D)\in\{0,\ast\}.

Proof.

Assume P⁡(h1,h2)=0P(h_{1},h_{2})=0. Then P⁡(h1,h2|D)=PD​(h1,h2)​P​(h1,h2)=PD​(h1,h2)⋅0P(h_{1},h_{2}|D)=P_{D}(h_{1},h_{2})P(h_{1},h_{2})=P_{D}(h_{1},h_{2})\cdot 0. This final term would normally simplify to 0, but will be ∗\ast if PD​(h1,h2)∈{∞,∗}P_{D}(h_{1},h_{2})\in\{\infty,\ast\}. ∎

If a Bayesian setup allows an outcome to be declared impossible with respect to another from one piece of evidence, and the reverse from another piece of evidence, those outcomes will become permanently incomparable. This failure mode will be familiar to those machine learning practitioners who have ever received a model full of NaNs.

Theorem 8.3.

If two outcomes are incomparable in a prior distribution, they will be incomparable in the posterior distribution. In other words, if P(h1,h2)=∗P(h_{1},h_{2})=\ast, then P(h1,h2|D)=∗P(h_{1},h_{2}|D)=\ast.

Proof.

Assume P(h1,h2)=∗P(h_{1},h_{2})=\ast. Then P(h1,h2|D)=PD(h1,h2)P(h1,h2)=PD(h1,h2)⋅∗=∗P(h_{1},h_{2}|D)=P_{D}(h_{1},h_{2})P(h_{1},h_{2})=P_{D}(h_{1},h_{2})\cdot\ast=\ast ∎

8.2 Example: A Noisy Channel

Here is an example of how relative probability gives us an interesting way of looking at statistical inference problems.

Suppose we are to receive a message in the outcome space Ω={0,1,…,K−1}\Omega=\{0,1,...,K-1\}. There is a probability of pp that the message goes through correctly. Otherwise, it gets scrambled and we receive a value in Ω\Omega drawn from the uniform distribution99 9 We could still have gotten lucky and received the correct value.. We receive the same message several times for redundancy, and we count ckc_{k} as the number of times the message was received as kk.

The indicator function can be used to get the absolute probability of receiving h1h_{1} given that the real message was h2h_{2}.

P(receivedh1|messageh2)=p[h1=h2]+1−pKP(\text{received}\>h_{1}|\text{message}\>h_{2})=p[h_{1}=h_{2}]+\frac{1-p}{K}

We then use this to construct an RPF for the likelihood ratio if we receive a single message, kk.

Pk​(h1,h2)=p[h1=k]+1−pKp[h2=k]+1−pK=pK[h1=k]+1−ppK[h2=k]+1−pP_{k}(h_{1},h_{2})=\frac{p[h_{1}=k]+\frac{1-p}{K}}{p[h_{2}=k]+\frac{1-p}{K}}=\frac{pK[h_{1}=k]+1-p}{pK[h_{2}=k]+1-p}

If we receive multiple messages in the count vector cc, we get the following likelihood formula:

Pc​(h1,h2)=∏k∈Ω(pK[h1=k]+1−ppK[h2=k]+1−p)ckP_{c}(h_{1},h_{2})=\prod_{k\in\Omega}\left(\frac{pK[h_{1}=k]+1-p}{pK[h_{2}=k]+1-p}\right)^{c_{k}}

Every factor of the product where k∉{h1,h2}k\notin\{h_{1},h_{2}\} will be 1 and have no effect. Therefore we can remove all those terms from the product.

Pc​(h1,h2)\displaystyle P_{c}(h_{1},h_{2}) =(pK[h1=h1]+1−ppK[h2=h1]+1−p)ch1​(pK[h1=h2]+1−ppK[h2=h2]+1−p)ch2\displaystyle=\left(\frac{pK[h_{1}=h_{1}]+1-p}{pK[h_{2}=h_{1}]+1-p}\right)^{c_{h_{1}}}\left(\frac{pK[h_{1}=h_{2}]+1-p}{pK[h_{2}=h_{2}]+1-p}\right)^{c_{h_{2}}} (11)
=(p​K+1−p1−p)ch1​(1−pp​K+1−p)ch2=(1+p​K1−p)ch1−ch2\displaystyle=\left(\frac{pK+1-p}{1-p}\right)^{c_{h_{1}}}\left(\frac{1-p}{pK+1-p}\right)^{c_{h_{2}}}=\left(1+\frac{pK}{1-p}\right)^{c_{h_{1}}-c_{h_{2}}}

Because the prior is uniform, the posterior is just equal to the likelihood.

P⁡(h1,h2|c)=Pc​(h1,h2)⋅P⁡(h1,h2)=(1+p​K1−p)ch1−ch2P(h_{1},h_{2}|c)=P_{c}(h_{1},h_{2})\cdot P(h_{1},h_{2})=\left(1+\frac{pK}{1-p}\right)^{c_{h_{1}}-c_{h_{2}}}

We now have an insight: the relative probability between two hypotheses is exponential on the difference between their counts. Formulating these problems in terms of relative probability often leads to easily interpretable results, even before converting into absolute probability (which may or may not be required). Using a different prior would be as easy as appending an additional term to the formula for P⁡(h1,h2|c)P(h_{1},h_{2}|c).

8.3 Digital Representation

Even if an inference problem starts with a mutually possible prior, the posterior could end up as any RPF including all types in figure 2. How can we represent an RPF digitally in a data structure that could account for all of these various possibilities?

Mutual possibility classes from section 5.4 provide a good framework for organizing an RPF. Every outcome is a member of a mutual possibility class, and has a relative probability within that class.

To that end, we need to maintain a mapping from outcomes to classes, and a partial order of all the mutual possibility classes. The data structure for this purpose could be as simple as listing the outbound edges for each class (illustrated by the graphs in figures 4, 5, and 6). There needs to be a method to compare two classes, and the comparison will return one of four values: greater, lesser, equal, or incomparable - or in numerical terms {∞,0,1,∗}\{\infty,0,1,\ast\}.

Each comparison might require a traversal of the mutual possibility graph which could get expensive in certain situations. The choice of data structure for partial orders comes with tradeoffs. For small outcome spaces graph traversal will usually be adequate. When Bayesian inference is performed, the mutual possibility classes will sometimes need to be split up - either vertically as some items in a class become impossible with respect to others - or horizontally as they become incomparable. The data structure should be able to account for this as well.

Let Ω\Omega be the outcome space, and let CC be the set of mutually possible classes on Ω\Omega with respect to the RPF. We maintain three functions

  • •

    α:Ω→C\alpha:\Omega\rightarrow C assigns outcomes to possibility classes.

  • •

    ℓ:Ω→(−∞,∞)\ell:\Omega\rightarrow(-\infty,\infty) provides a log value for an outcome within its mutual possibility class. This allows us to use the full range of floating point numbers available on our machine.

  • •

    Q:C×C→{∞,0,1,∗}Q:C\times C\rightarrow\{\infty,0,1,\ast\} compares two probability classes.

From these, we can compute the value of an RPF with the following formula:

P⁡(h1,h2)=Q⁡(α⁡(h1),α⁡(h2))⋅eℓ⁡(h1)−ℓ⁡(h2)P(h_{1},h_{2})=Q(\alpha(h_{1}),\alpha(h_{2}))\cdot e^{\ell(h_{1})-\ell(h_{2})}

The simplest RPF to represent in this form is the uniform RPF. In this case, there is a single possibility class so C={0}C=\{0\}, and ℓ\ell is constant so it can be set as ℓ⁡(h)=0\ell(h)=0. No data is needed for Q because it has an identity rule where Q⁡(0,0)=1Q(0,0)=1. In fact, Q is itself an RPF on C where instead of providing values in 𝕄∗\mathbb{M}^{\ast} it provides values in the subspace {∞,0,1,∗}\{\infty,0,1,\ast\}.

9 Topology and Limits in Relative Probability Space

One of the benefits of relative probability spaces is their properties with respect to limits. This section will show that limits of totally comparable RPFs are also totally comparable RPFs. In other words, the space of RPFs is topologically closed.

This effort caps off a significant argument in favor of relative probability. RPFs hold on to certain pieces information under the limit operation, while absolute probability does not. Some background in topology1010 10 Mendelson (1990) [12] and Bradley et al. (2020) [13] the formal definitions and theorems which we will work off of here. required for this section.

9.1 Relative Probability Spaces

Definition 9.1.

RPF∗​(K)\text{RPF}^{\ast}(K) is the set of relative probability functions of size K (where Ω={0,1,…,K−1}\Omega=\{0,1,...,K-1\}). Likewise RPF​(K)\text{RPF}(K) is the set of all totally comparable RPFs of size K.

Because the set of absolute distributions with |Ω|=K|\Omega|=K is embedded in ℝK\mathbb{R}^{K} as seen in figure 1, its topological properties are well understood. The simplex is closed, bounded, and compact.

For relative probability distributions, there is no obvious way to embed it into K-dimensional euclidean space1111 11 Though it should be possible! See section 10.3.. The relative probability space is more complicated, because at the corners and edges of the simplex lurk entire subspaces where zero-probability outcomes are still being compared in different configurations.

Fortunately, RPF​(K)\text{RPF}(K) can still be embedded into a much larger euclidean space. Any P∈RPF​(K)P\in\text{RPF}(K) is a function of type Ω×Ω→𝕄\Omega\times\Omega\rightarrow\mathbb{M} that satisfies the fundamental properties. Therefore RPF​(K)\text{RPF}(K) can at least be embedded into 𝕄K2\mathbb{M}^{K^{2}}.

A topology can be defined by a basis of open sets, and for euclidean space (and metric spaces more generally) the basis is the set of open balls on ℝn\mathbb{R}^{n}.

Definition 9.2.

An open ball of size ϵ\epsilon around point x∈ℝnx\in\mathbb{R}^{n} is the set of all points y such that |x−y|<ϵ|x-y|<\epsilon.

Open balls do not work as a basis on the set 𝕄\mathbb{M} because an open ball around ∞\infty would contain only ∞\infty when we instead want them to also contain some interval (x,∞](x,\infty]. This is remedied by using the following transformation between 𝕄\mathbb{M} and [0,1][0,1]. We take this transformation to be continuous, or topology conserving, so that the topology of 𝕄\mathbb{M} is defined by the topology of [0,1][0,1].

Definition 9.3.

The inverse odds transform is the function odds−1:𝕄→[0,1]\text{odds}^{-1}:\mathbb{M}\rightarrow[0,1] with odds−1​(0)=0\text{odds}^{-1}(0)=0 and odds−1​(∞)=1\text{odds}^{-1}(\infty)=1 defined by

odds−1​(x)=xx+1\text{odds}^{-1}(x)=\frac{x}{x+1}

The inverse odds transformation establishes 𝕄\mathbb{M} as topologically equivalent to the closed interval [0,1][0,1]. Now RPF​(K)\text{RPF}(K) can be embedded into a bounded region of euclidean space, namely [0,1]K2[0,1]^{K^{2}} by applying inverse odds to it. The result won’t follow the fundamental axioms, but it will respect all of the topological properties. Appendix A provides an alternative way to define a topology on RPF​(K)\text{RPF}(K).

9.2 Limit Points and Compactness

We now present an argument for the closure of RPF​(K)\text{RPF}(K) under limits.

Definition 9.4.

xx is a limit point of set AA if any open set containing xx also contains points in AA. Equivalently for euclidean space, any open ball containing xx also intersects with AA.

Theorem 9.1.

Let PP be a limit point of RPF​(K)\text{RPF}(K). Then PP satisfies all of the fundamental axioms and is therefore a member of RPF​(K)\text{RPF}(K). In other words, RPF​(K)\text{RPF}(K) is topologically closed.

Proof.

Let PP be a limit point of RPF​(K)\text{RPF}(K). We can show that for each fundamental axiom, if PP doesn’t satisfy the axiom then some open ball around PP also doesn’t satisfy the axiom.

Identity: If PP doesn’t satisfy the identity axiom, then P⁡(h,h)≠1P(h,h)\neq 1 for some outcome hh. The inverse odds transform maps the value 11 to 12\frac{1}{2}, so a choice of ϵ\epsilon less than |odds−1​(P⁡(h,h))−12||\text{odds}^{-1}(P(h,h))-\frac{1}{2}| will contain only elements P′P^{\prime} where P′​(h,h)≠1P^{\prime}(h,h)\neq 1.

Inverse: A similar argument applies here if for some pair h1h_{1} and h2h_{2}, P⁡(h1,h2)≠P⁡(h1,h2)P(h_{1},h_{2})\neq P(h_{1},h_{2}). An ϵ\epsilon can be selected that is small enough so that P⁡(h1,h2)P(h_{1},h_{2}) is never equal to P⁡(h1,h2)P(h_{1},h_{2})

Composition: Suppose P⁡(a,c):≇P⁡(a,b)⋅P⁡(b,c)P(a,c):\ncong P(a,b)\cdot P(b,c). This can only be true if P⁡(a,c)≠P⁡(a,b)⋅P⁡(b,c)P(a,c)\neq P(a,b)\cdot P(b,c). But then this gap allows us to make the same argument from before; for some sufficiently small ϵ\epsilon-ball around PP, the composition axiom will still be false.

This means that only functions PP which satisfy the fundamental axioms can be limit points to RPF​(K)\text{RPF}(K), and therefore RPF​(K)\text{RPF}(K) contains all its limit points and is closed. ∎

We can now prove that RPF​(K)\text{RPF}(K) is compact thanks to the Heine-Borel theorem, stated as follows (wording from Bradley [13]).

Theorem 9.2 (Heine-Borel Theorem).

A subset of ℝn\mathbb{R}^{n} is compact if and only if it is closed and bounded.

Theorem 9.3.

RPF​(K)\text{RPF}(K) is compact.

Proof.

Theorem 9.1 states that RPF​(K)\text{RPF}(K) is closed, and by the odds transform it is also bounded. By the Heine-Borel theorem, it is compact. ∎

10 Future Work

10.1 Expansions to infinite spaces

The obvious extension to this work is to expand relative probability to a generalized space which may be infinite, and thus capture all of the variety of probability distributions that one might wish to define. Section 6 provides a good framework for this which would start by modifying statement 3 to ask for an additive property. The relative probability function would then become a relative probability distribution.

This process raises certain questions.

  1. 1.

    Relative probability would provide an interesting basis for the theory of finitely additive probability measures, or charges[14]. If we do not require countable additivity and instead opt for the weaker finite additivity, this would allow for constructions like a fair countable lottery (also known as a De Finetti Lottery[15]). In the interest of studying limits of RPFs, this would be desirable. Finite additivity may also be a better mathematical model for computational methods in probability, on which additions must always be finite.

  2. 2.

    If we derive a notion of probability density, then can these densities at a particular pair of events be used to compare the relative probability of those events? What specific properties of the relative probability distribution are required to make this work? Can we start with a density function and work upwards?

It appears possible to use these ideas to create a unified version of the Hausdorff measure - which finds the size of an object given its dimension. Instead of considering it to be multiple measures, we can have a single measure where bounded sets of equal dimension are mutually possible, and smaller-dimensional objects are mutually impossible with respect to larger dimensional objects.

10.2 Relationship to Category Theory

Category theorists will recognize that an RPF describes a thin category where any pair of objects have at most one morphism connecting them (per direction). The relative probability axioms can be analyzed and approached through the lens of category theory in order to learn more about them.

CCBBAA63216\frac{1}{6}13\frac{1}{3}12\frac{1}{2}111

The recent work of Censi et al.[16] concerns negative information in categories, which corresponds to the wildcard element ∗\ast. It represents regions of the probability function that remain incomparable. This work could be used to subsume, develop, and refine the indeterminate wildcard concept.

10.3 Embedding in Lower Dimensional Euclidean Space

Absolute probability functions have this advantage where they can be embedded into a simplex in ℝK\mathbb{R}^{K}. For relative probability functions it is not so straightforward but it should still be possible given that RPF​(K)\text{RPF}(K) is a (K-1)-dimensional set that can be embedded into [0,1]K2[0,1]^{K^{2}}. For example, the space RPF​(3)\text{RPF}(3) can be mapped as a hexagon, where each point can be assigned a probability based on its distance between two parallel sides as in figure 8.

P⁡(a)=1P(a)=1P⁡(c)=0P(c)=0P⁡(b)=1P(b)=1P⁡(a)=0P(a)=0P⁡(c)=1P(c)=1P⁡(b)=0P(b)=0
Figure 8: A triangle probability simplex embedded into ℝ2\mathbb{R}^{2} as a hexagon.

In this case, the triangular simplex has been truncated. For higher order simplices, this appears to become increasingly unwieldy unless some simplifying trick is developed. If it is successfully done, then a cleaner representation for members of RPF​(K)\text{RPF}(K) becomes available.

11 Conclusion

Modeling probability as a “relative” measure takes on a new significance given the proliferation of computational methods. The nice properties of RPFs include formula simplification, closure under limits, and the ability to model non-standard distributions.

Should practitioners regularly use the language of relative probability when describing and deploying probabilistic models? It may not be beneficial in every situation, but they should certainly consider it! Computational techniques for Bayesian inference and machine learning in many cases remove the need to normalize the distributions involved, so why not make it official?

Appendix A Topology and Limits in Relative Probability Space

In section 9, we established a topology on RPF​(K)\text{RPF}(K) by embedding it into Euclidean space, and demonstrated its compactness. Here is an alternative topology that defines a basis of open sets directly. While this is not used in the context of this paper, it shows promise in providing a more natural framework for topological proofs and simplex partitioning.

The notion of an open set changes when a topological space is restricted to a lower dimension. For example, on the real number line ℝ\mathbb{R}, the open interval (0, 1) is an open set. However, once it is embedded into ℝ2\mathbb{R}^{2}, it is now a line segment in a plane and no longer open (see figure 9). For example, the set {(x,y):x∈(0,1)​and​y∈(−ϵ,+ϵ)}\{(x,y):x\in(0,1)\;\text{and}\;y\in(-\epsilon,+\epsilon)\} given an ϵ>0\epsilon>0 is such an open set on ℝ2\mathbb{R}^{2}. The line segment (0, 1) is open in ℝ\mathbb{R} because it is the restriction of this open set on ℝ2\mathbb{R}^{2}.

xxyy01−ε-\varepsilonε\varepsilon
Figure 9: The small box that is the interior of the dotted rectangle is an open set in ℝ2\mathbb{R}^{2}, and therefore its restriction to ℝ\mathbb{R} - the line segment - is an open set in ℝ\mathbb{R}. But the line segment is not open in ℝ2\mathbb{R}^{2}.

Likewise, an open set on RPF​(K)\text{RPF}(K) will not be an open set on RPF​(K+1)\text{RPF}(K+1). Starting with RPF​(2)\text{RPF}(2), we find a totally comparable RPF that corresponds 1:1 with the magnitude space.

Theorem A.1.

Let Ω={0,1}\Omega=\{0,1\} have two elements, with relative probability function PP. Then, PP is completely determined by P⁡(0,1)P(0,1).

Proof.

Let q=P⁡(0,1)q=P(0,1). By the inverse symmetric property, P⁡(1,0)=q−1P(1,0)=q^{-1}. These values completely determine PP on the outcome level. ∎

This gives us both a topology and a compactness proof for RPF​(2)\text{RPF}(2) because it is isomorphic to 𝕄\mathbb{M}. Its basis for open sets are the open intervals of 𝕄\mathbb{M}, including those intervals that include 0 and ∞\infty. For K>2K>2, we will need more powerful tools, namely open patches. The set of open patches, which come in several flavors, will be a basis for the open sets on RPF​(K)\text{RPF}(K).

Definition A.1.

An interior open patch of RPF​(K)\text{RPF}(K) is one of the following:

  1. 1.

    If K=2K=2, a subset parameterized by an interior open interval of magnitudes. {P|a<P⁡(h1,h2)<b}\{P|a<P(h_{1},h_{2})<b\} for some a,b∈𝕄a,b\in\mathbb{M}

  2. 2.

    If K>2K>2, a composition of interior patches with composing function P⊤P_{\top} also being an interior patch.

Interior open patches contain only totally mutually possible functions as illustrated in figure 10.

P⊤P_{\top}P0P_{0}ABC
)(Interior OpenPatchAABBCC)(PTP_{T}P0P_{0})(
Figure 10: An interior open patch captures a contiguous set inside the probability simplex. Above is an example of a composite probability distribution where the diagram on the left shows how P⊤P_{\top} and P0P_{0} are composed, and the right shows how the interior segments on both interact to form a patch.
Definition A.2.

A facet1212 12 A facet of a simplex is a subset where one parameter is equal to zero - equivalent to a face on a 3D object. patch of RPF​(K)\text{RPF}(K) is one of the following:

  1. 1.

    If K=2K=2, an interval of the form {P|0<P⁡(h1,h2)<a}\{P|0<P(h_{1},h_{2})<a\} for some a∈𝕄a\in\mathbb{M}

  2. 2.

    If K>2K>2, a composition where P⊤P_{\top} is drawn from an interior open patch, and all but one of the components are drawn from interior open patches. The final component - the facet component - is itself drawn from a facet patch.

Definition A.3.

An exterior open patch is a one of the following:

  1. 1.

    A facet patch.

  2. 2.

    A composition where P⊤P_{\top} is a facet patch. The facet component is itself drawn from any open patch, and all the other components are drawn from interior open patches.

As seen in figure 11, exterior open patches touch the hyperfaces (facets) of the simplex as well as the vertices and edges. As the number of dimensions increases and the composition diagram changes, more permutations are possible.

)AABBCC)(PTP_{T}P0P_{0})
)AABBCC(PTP_{T}P0P_{0})
AABBCC()(PTP_{T}P0P_{0}P0P_{0})(
Figure 11: Exterior open patches. On the left is the facet patch, because it only touches a side (facet) of the simplex and not a corner. In the center is an exterior open patch where the facet component P0P_{0} is itself a facet patch (touching an edge and a corner), and on the right is an exterior open patch that touches a corner only because P0P_{0} is an interior open patch. Note that the point containing the corner at CC in the middle and right diagram is only half filled because the patch contains some values where P⁡(C)=1P(C)=1 and not others, depending on the relative probability between AA and BB.
Definition A.4.

An open patch is a subset of RPF​(K)\text{RPF}(K) that is either an interior or exterior open patch.

Now let the open patches be a basis for an open set thus defining a topology on RPF​(K)\text{RPF}(K).

Definition A.5.

An open set of RPF​(K)\text{RPF}(K) is any (potentially infinite) union of open patches on RPF​(K)\text{RPF}(K), or any finite intersection of open patches on RPF​(K)\text{RPF}(K).

These open patches are useful building blocks to building recursive proofs of compactness and other topological properties we might want to prove in the future.

References

  • [1] A. N. Kolmogorov. Foundations of the Theory of Probability. Chelsea Publishing Company, New York (1956).
  • [2] Lyon, A. (2016). Kolmogorov’s Axiomatisation and its Discontents. The Oxford handbook of probability and philosophy, 155-166.
  • [3] November, D. D. (2018). Zero Probability.
  • [4] Hájek, A. (2003). What conditional probability could not be. Synthese, 137(3), 273-323.
  • [5] Rényi, A. (1955). On a new axiomatic theory of probability. Acta Mathematica Hungarica, 6(3-4), 285-335.
  • [6] Kohlberg, E., & Reny, P. J. (1997). Independence on relative probability spaces and consistent assessments in game trees. Journal of Economic Theory, 75(2), 280-313.
  • [7] Heinemann, F. (1997). Relative Probabilities. Working paper, http://www. sfm. vwl. uni-muenchen. de/heinemann/publics/relative probabilities-intro. htm.
  • [8] Heinemann, F. (1995). Closing Economic Models by Information Assumptions. Inst. für Volkswirtschaftslehre u. Statistik d. Univ. Mannheim.
  • [9] Matoušek, J., & Nešetřil, J. (2008). Invitation to discrete mathematics. OUP Oxford.
  • [10] Kahan, W. (1996). IEEE standard 754 for binary floating-point arithmetic. Lecture Notes on the Status of IEEE, 754(94720-1776), 11.
  • [11] Sklar, M. (2022). Sampling Bias Correction for Supervised Machine Learning: A Bayesian Inference Approach with Practical Applications. arXiv preprint arXiv:2203.06239.
  • [12] Mendelson, B. (1990). Introduction to topology. Courier Corporation.
  • [13] Bradley, T. D., Bryson, T., & Terilla, J. (2020). Topology: A Categorical Approach. MIT Press.
  • [14] Rao, K. B., & Rao, M. B. (1983). Theory of charges: a study of finitely additive measures. Academic Press.
  • [15] Finetti, B. D. (1974). Theory of probability, vol. 1 and vol. 2.
  • [16] Censi, A., Frazzoli, E., Lorand, J., & Zardini, G. (2022). Categorification of Negative Information using Enrichment. arXiv preprint arXiv:2207.13589.

Funding and Interests

Local Maximum Labs is an ongoing effort to create and disseminate knowledge on intelligent computing, and its current resource is only the author’s time. No funds, grants, or other support was received to assist with the preparation of this manuscript. The author certifies that they have no affiliations with or involvement in any external organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.